I don't understand this concept. Please help.
Write the equation of the line that passes through each pair of points in slope-intercept form.
1. (5, -1) and (-7, -4)
2. (5,1) and (-1, -5)
3. (4,9) and (-5,3)

Answer by checkley77(12844)

Y=mX+b where m=slope & b=the Y intercept.
Slope=(Y2-Y1)/(X2-X1)
1. (5, -1) and (-7, -4)
Slope=(-4+1)/(-7-5)
Slope=-3/-12
Slope=1/4
-1=1/4*5+b
-1=5/4+b
b=-1-5/4
b=-4/4-5/4
b=-9/4 The Y intercept.
Y=x/4-9/4 ans.
2. (5,1) and (-1, -5)
Slope=(-5-1)/(-1-5)
Slope=-6/-6
Slope=1
1=1*5+b
1=5+b
b=1-5
b=-4 The Y intercept.
Y=X-4 ans.
3. (4,9) and (-5,3)
Slope=(3-9)/(-5-4)
Slope=-6/-9
Slope=2/3
9=2/3*4+b
9=8/3+b
b=9-8/3
b=27/3-8/3
b=19/3
Y=2x/3+19/3 ans.